ω=120rev/min then it changed to 180rev/min in∆θ=4rev

first find angular acceleration and then find ∆t


1
Expert's answer
2021-06-18T11:23:41-0400
"∆θ=\omega_it+0.5\alpha t^2\\2∆θ=(\omega_i+\omega_f)t
\\2(4)=(120+180)t\\t=1.6\ s\\
4(2\pi)=(1.6)(120)\frac{2\pi}{60}+0.5(1.6)^2\alpha\\
\alpha=3.9\frac{rad}{s^2}"